# MHNet: Multi-view High-order Network for Diagnosing Neurodevelopmental Disorders Using Resting-state fMRI

### v1

@@ 5 Results
5.1 Classification Performance
@@
Table 2: Classification performance for ASD (ASD vs. NC) on ABIDE I and ABIDE II, and for ADHD (ADHD vs. NC) on ADHD-200, using the AAL1 atlas. Results (shown as percentages) were derived from 10-fold cross-validation. The best result in each category is highlighted in bold.
@@
ABIDE-I SVM 61.46 ± 2.30 60.51 ± 2.81 60.60 ± 5.24 59.96 ± 1.90 60.13 ± 2.56
KNN 60.40 ± 3.05 58.68 ± 3.48 58.40 ± 1.51 60.49 ± 2.23 63.78 ± 2.12
MLP 62.06 ± 5.13 61.61 ± 4.54 62.93 ± 4.49 59.57 ± 3.07 63.44 ± 3.08
GCN 64.14 ± 5.39 63.49 ± 1.06 64.10 ± 1.35 60.73 ± 5.99 61.74 ± 0.28
BrainGNN 67.70 ± 2.15 65.70 ± 1.40 69.79 ± 0.51 65.73 ± 1.12 67.23 ± 2.53
AL-NEGAT 71.54 ± 2.58 69.33 ± 2.60 71.94 ± 1.11 68.63 ± 2.59 70.86 ± 4.22
MVS-GCN 69.30 ± 1.65 67.39 ± 3.36 67.64 ± 1.46 66.59 ± 4.41 67.44 ± 0.84
Hi-GCN 71.83 ± 2.41 68.60 ± 2.86 68.62 ± 1.28 65.84 ± 3.31 68.00 ± 2.30
BrainNetCNN 68.24 ± 3.37 69.10 ± 0.33 72.38 ± 1.96 69.67 ± 2.51 69.45 ± 2.20
BrainNetTF 71.13 ± 2.57 70.15 ± 2.21 71.09 ± 1.46 69.25 ± 1.99 71.54 ± 2.91
MAHGCN 72.56 ± 2.88 72.06 ± 2.53 72.77 ± 2.66 70.79 ± 0.85 71.44 ± 3.85
Com-BrainTF 73.06 ± 1.45 71.81 ± 2.80 73.12 ± 0.72 69.14 ± 1.37 71.73 ± 3.91
@@
ABIDE-II SVM 60.59 ± 1.67 59.62 ± 1.29 61.23 ± 4.80 58.54 ± 3.24 60.00 ± 4.39
KNN 61.81 ± 1.75 60.38 ± 2.51 63.29 ± 2.83 59.63 ± 5.61 62.54 ± 3.12
MLP 63.35 ± 3.27 60.27 ± 3.65 61.27 ± 1.59 59.03 ± 4.36 61.84 ± 2.98
GCN 63.96 ± 2.53 63.22 ± 2.39 64.54 ± 2.06 60.13 ± 2.59 63.27 ± 3.90
BrainGNN 68.70 ± 1.09 68.35 ± 1.40 69.57 ± 0.51 67.73 ± 1.12 69.03 ± 3.53
AL-NEGAT 71.63 ± 1.26 70.87 ± 3.03 71.03 ± 2.43 68.54 ± 1.48 70.32 ± 3.22
MVS-GCN 70.52 ± 1.65 68.69 ± 2.16 69.34 ± 2.76 67.18 ± 2.41 69.35 ± 2.39
Hi-GCN 70.04 ± 2.41 70.37 ± 1.23 71.24 ± 2.56 68.23 ± 2.01 71.24 ± 2.30
BrainNetCNN 69.24 ± 1.29 68.10 ± 1.73 71.38 ± 1.96 67.67 ± 1.35 70.45 ± 2.20
BrainNetTF 72.13 ± 2.57 69.15 ± 2.21 71.09 ± 1.46 67.25 ± 1.99 69.54 ± 2.91
MAHGCN 71.39 ± 1.46 71.06 ± 0.94 70.77 ± 1.42 70.47 ± 0.97 71.64 ± 1.39
Com-BrainTF 71.89 ± 0.75 70.81 ± 1.57 72.02 ± 0.94 69.14 ± 1.62 70.39 ± 3.36
@@
ADHD-200 SVM 60.60 ± 2.88 59.76 ± 2.68 59.61 ± 4.14 59.34 ± 2.50 59.02 ± 2.77
KNN 58.46 ± 2.62 58.01 ± 2.23 57.98 ± 1.34 59.73 ± 1.34 60.81 ± 1.09
MLP 60.99 ± 4.04 60.65 ± 4.21 62.10 ± 2.96 58.93 ± 2.65 62.20 ± 3.19
GCN 63.46 ± 2.62 65.78 ± 1.09 63.61 ± 1.88 62.72 ± 5.25 64.29 ± 0.95
BrainGNN 64.07 ± 2.81 67.16 ± 1.50 61.93 ± 0.49 60.73 ± 1.51 62.34 ± 2.15
AL-NEGAT 69.33 ± 3.56 68.60 ± 2.91 65.27 ± 1.97 63.14 ± 2.52 66.88 ± 3.52
MVS-GCN 67.49 ± 2.91 66.72 ± 3.05 68.80 ± 1.60 65.11 ± 4.13 67.53 ± 1.11
Hi-GCN 68.84 ± 2.15 67.39 ± 2.76 69.42 ± 2.74 65.39 ± 3.33 67.15 ± 3.90
BrainNetCNN 68.33 ± 2.32 66.71 ± 1.30 64.81 ± 1.48 64.01 ± 1.41 66.38 ± 2.90
BrainNetTF 69.26 ± 3.09 67.64 ± 2.39 69.53 ± 0.83 67.70 ± 2.92 68.06 ± 1.53
MAHGCN 68.34 ± 3.56 68.13 ± 1.88 68.22 ± 2.67 67.91 ± 1.06 68.67 ± 3.34
Com-BrainTF 70.10 ± 2.65 67.73 ± 2.45 68.22 ± 0.96 69.59 ± 0.73 69.23 ± 3.75
@@
Table 3: Classification performance for ASD (ASD vs. NC) on ABIDE I and ABIDE II, and for ADHD (ADHD vs. NC) on ADHD-200, using the Brainnetome atlas. Results (shown as percentages) were derived from 10-fold cross-validation. The best result in each category is highlighted in bold.
@@
ABIDE-I SVM 63.08 ± 3.19 62.37 ± 3.82 62.30 ± 4.83 61.57 ± 2.84 61.57 ± 3.35
KNN 61.16 ± 3.39 60.14 ± 2.65 60.51 ± 2.56 62.71 ± 1.35 63.16 ± 1.17
MLP 64.03 ± 3.89 63.20 ± 4.62 64.58 ± 3.25 61.73 ± 2.80 64.37 ± 2.46
GCN 65.10 ± 6.59 64.35 ± 1.62 66.13 ± 2.34 59.68 ± 6.52 63.82 ± 1.27
BrainGNN 69.94 ± 2.62 68.68 ± 1.62 66.28 ± 0.62 65.68 ± 0.62 66.23 ± 2.36
AL-NEGAT 71.29 ± 3.19 70.37 ± 2.54 71.13 ± 2.17 68.21 ± 2.39 70.29 ± 3.10
MVS-GCN 72.51 ± 2.12 71.33 ± 2.92 72.13 ± 2.42 69.24 ± 3.85 71.30 ± 1.39
Hi-GCN 74.55 ± 3.23 72.31 ± 2.64 70.34 ± 2.53 68.03 ± 2.45 71.81 ± 3.59
BrainNetCNN 72.59 ± 2.91 71.56 ± 0.93 72.21 ± 1.35 67.59 ± 1.31 72.24 ± 2.85
BrainNetTF 74.93 ± 2.37 73.28 ± 1.97 72.11 ± 1.27 68.53 ± 2.71 70.89 ± 1.73
MAHGCN 73.21 ± 3.02 73.52 ± 1.63 73.02 ± 2.14 71.55 ± 1.38 72.57 ± 3.03
Com-BrainTF 74.12 ± 1.72 73.58 ± 1.81 74.31 ± 0.51 70.64 ± 0.38 74.91 ± 3.41
@@
ABIDE-II SVM 61.43 ± 2.23 60.15 ± 2.37 61.25 ± 3.74 59.44 ± 1.06 60.57 ± 4.25
KNN 62.37 ± 1.09 61.24 ± 1.76 63.83 ± 1.67 60.82 ± 2.15 62.07± 2.37
MLP 62.83 ± 2.42 62.02 ± 3.00 63.15 ± 2.61 60.54 ± 2.41 63.15 ± 1.98
GCN 63.80 ± 2.76 62.85 ± 2.03 64.96 ± 3.62 58.17 ± 3.87 62.65 ± 1.84
BrainGNN 69.37 ± 3.27 66.34 ± 2.35 67.71 ± 2.49 65.28 ± 3.74 67.77 ± 2.75
AL-NEGAT 70.14 ± 0.80 68.11 ± 0.57 70.95 ± 3.41 68.95 ± 3.22 69.11 ± 3.55
MVS-GCN 69.24 ± 3.93 67.01 ± 2.30 69.97 ± 2.12 67.80 ± 3.23 68.92 ± 3.91
Hi-GCN 70.30 ± 2.74 69.94 ± 1.00 70.17 ± 3.81 68.64 ± 2.33 69.86 ± 1.95
BrainNetCNN 69.32 ± 1.43 67.13 ± 3.21 68.02 ± 2.10 65.14 ± 2.49 66.98 ± 0.57
BrainNetTF 71.42 ± 2.66 70.89 ± 2.64 71.97 ± 1.66 69.03 ± 2.80 71.76 ± 2.89
MAHGCN 72.17 ± 1.76 71.27 ± 2.03 72.56 ± 2.94 69.89 ± 0.71 71.33 ± 2.83
Com-BrainTF 72.73 ± 2.85 71.18 ± 1.24 72.92 ± 0.95 69.41 ± 1.60 71.35 ± 1.77
@@
ADHD-200 SVM 61.15 ± 2.21 60.20 ± 3.00 60.30 ± 4.31 59.65 ± 2.04 59.83 ± 3.14
KNN 59.09 ± 2.71 58.38 ± 2.42 58.09 ± 1.61 60.18 ± 0.78 61.47 ± 0.70
MLP 61.75 ± 3.55 61.30 ± 3.82 62.63 ± 2.74 59.26 ± 2.19 63.14 ± 2.34
GCN 60.84 ± 5.93 63.19 ± 0.97 63.79 ± 1.68 57.42 ± 5.57 61.43 ± 3.56
BrainGNN 64.66 ± 2.62 62.78 ± 1.62 62.33 ± 0.62 61.21 ± 0.62 63.00 ± 2.36
AL-NEGAT 64.03 ± 3.19 63.21 ± 2.54 62.79 ± 2.17 63.95 ± 2.39 63.90 ± 3.10
MVS-GCN 66.96 ± 2.12 64.04 ± 2.92 66.33 ± 2.42 63.98 ± 3.85 65.05 ± 1.39
Hi-GCN 67.49 ± 3.23 66.97 ± 2.64 65.84 ± 2.53 64.14 ± 2.45 66.08 ± 3.59
BrainNetCNN 69.08 ± 2.91 68.31 ± 0.93 69.42 ± 1.35 65.64 ± 1.31 68.08 ± 2.85
BrainNetTF 68.90 ± 2.37 68.26 ± 1.97 68.29 ± 1.27 66.03 ± 2.71 67.37 ± 1.73
MAHGCN 69.56 ± 3.02 67.76 ± 1.63 68.88 ± 2.14 68.33 ± 1.38 68.47 ± 3.03
Com-BrainTF 70.12 ± 1.72 68.14 ± 1.81 70.99 ± 0.51 67.05 ± 0.38 69.25 ± 3.41
@@ 5.2.1 Influence of Cutoff Threshold
For HGNN, we use the obtained RV coefficient matrix to construct the adjacency matrix of the brain hierarchical network of each data set, but the RV coefficient matrix is not a sparse binary matrix. In order to achieve the best classification effect, threshold processing is required for the matrix. Specifically, if the RV coefficient between brain area $i$ and brain area $j$ is less than the cutoff threshold T, the value at $(i,j)$ of the RV coefficient matrix is set to 0, otherwise it is set to 1. As shown in Figure 2, in order to select a suitable cutoff threshold, we plotted the relationship curve between the percentage of retained edges and the cutoff threshold, and selected the inflection point of the curve as the cutoff point. The cutoff threshold at this point is used to process the RV coefficient matrix into a sparse binary adjacency matrix.

### v2

@@ 5 Results
5.1 Classification Performance and Model Efficiency
@@
The proposed MHNet maintains a balanced trade-off between model complexity and computational efficiency. The input dimension varies across datasets due to different ROI definitions (e.g., 116 ROIs for AAL1 atlas). The network contains approximately 0.63M trainable parameters, with the HGNN module accounting for about 64% of the total parameters due to its three parallel graph neural networks and the HCNN module accounting for about 36% of the total parameters. During training on a single NVIDIA RTX 3090 GPU, the model requires around 16 GB of GPU memory. The inference training times on ABIDE-I, ABIDE-II, ADHD-200 are about 4.7 hours, 2.8 hours, 3.5 hours respectively.
Table 2: Classification performance for ASD (ASD vs. NC) on ABIDE I and ABIDE II, and for ADHD (ADHD vs. NC) on ADHD-200, using the AAL1 atlas. Results (shown as percentages) were derived from 10-fold cross-validation. Statistically significant differences from MHNet marked *. The best result in each category is highlighted in bold.
@@
ABIDE-I SVM 61.46 ± 2.30* 60.51 ± 2.81* 60.60 ± 5.24* 59.96 ± 1.90* 60.13 ± 2.56*
KNN 60.40 ± 3.05* 58.68 ± 3.48* 58.40 ± 1.51* 60.49 ± 2.23* 63.78 ± 2.12*
MLP 62.06 ± 5.13* 61.61 ± 4.54* 62.93 ± 4.49* 59.57 ± 3.07* 63.44 ± 3.08*
GCN 64.14 ± 5.39* 63.49 ± 1.06* 64.10 ± 1.35* 60.73 ± 5.99* 61.74 ± 0.28*
BrainGNN 67.70 ± 2.15* 65.70 ± 1.40* 69.79 ± 0.51* 65.73 ± 1.12* 67.23 ± 2.53*
AL-NEGAT 71.54 ± 2.58* 69.33 ± 2.60* 71.94 ± 1.11* 68.63 ± 2.59* 70.86 ± 4.22*
MVS-GCN 69.30 ± 1.65* 67.39 ± 3.36* 67.64 ± 1.46* 66.59 ± 4.41* 67.44 ± 0.84*
Hi-GCN 71.83 ± 2.41* 68.60 ± 2.86* 68.62 ± 1.28* 65.84 ± 3.31* 68.00 ± 2.30*
BrainNetCNN 68.24 ± 3.37* 69.10 ± 0.33* 72.38 ± 1.96* 69.67 ± 2.51 69.45 ± 2.20*
BrainNetTF 71.13 ± 2.57* 70.15 ± 2.21* 71.09 ± 1.46* 69.25 ± 1.99* 71.54 ± 2.91*
MAHGCN 72.56 ± 2.88* 72.06 ± 2.53* 72.77 ± 2.66* 70.79 ± 0.85 71.44 ± 3.85
Com-BrainTF 73.06 ± 1.45* 71.81 ± 2.80* 73.12 ± 0.72* 69.14 ± 1.37* 71.73 ± 3.91*
@@
ABIDE-II SVM 60.59 ± 1.67* 59.62 ± 1.29* 61.23 ± 4.80* 58.54 ± 3.24* 60.00 ± 4.39*
KNN 61.81 ± 1.75* 60.38 ± 2.51* 63.29 ± 2.83* 59.63 ± 5.61* 62.54 ± 3.12*
MLP 63.35 ± 3.27* 60.27 ± 3.65* 61.27 ± 1.59* 59.03 ± 4.36* 61.84 ± 2.98*
GCN 63.96 ± 2.53* 63.22 ± 2.39* 64.54 ± 2.06* 60.13 ± 2.59* 63.27 ± 3.90*
BrainGNN 68.70 ± 1.09* 68.35 ± 1.40* 69.57 ± 0.51* 67.73 ± 1.12* 69.03 ± 3.53*
AL-NEGAT 71.63 ± 1.26* 70.87 ± 3.03* 71.03 ± 2.43* 68.54 ± 1.48* 70.32 ± 3.22*
MVS-GCN 70.52 ± 1.65* 68.69 ± 2.16* 69.34 ± 2.76* 67.18 ± 2.41* 69.35 ± 2.39*
Hi-GCN 70.04 ± 2.41* 70.37 ± 1.23* 71.24 ± 2.56* 68.23 ± 2.01* 71.24 ± 2.30*
BrainNetCNN 69.24 ± 1.29* 68.10 ± 1.73* 71.38 ± 1.96* 67.67 ± 1.35* 70.45 ± 2.20*
BrainNetTF 72.13 ± 2.57* 69.15 ± 2.21* 71.09 ± 1.46* 67.25 ± 1.99* 69.54 ± 2.91*
MAHGCN 71.39 ± 1.46 71.06 ± 0.94* 70.77 ± 1.42* 70.47 ± 0.97 71.64 ± 1.39*
Com-BrainTF 71.89 ± 0.75* 70.81 ± 1.57* 72.02 ± 0.94* 69.14 ± 1.62* 70.39 ± 3.36*
@@
ADHD-200 SVM 60.60 ± 2.88* 59.76 ± 2.68* 59.61 ± 4.14* 59.34 ± 2.50* 59.02 ± 2.77*
KNN 58.46 ± 2.62* 58.01 ± 2.23* 57.98 ± 1.34* 59.73 ± 1.34* 60.81 ± 1.09*
MLP 60.99 ± 4.04* 60.65 ± 4.21* 62.10 ± 2.96* 58.93 ± 2.65* 62.20 ± 3.19*
GCN 63.46 ± 2.62* 65.78 ± 1.09* 63.61 ± 1.88* 62.72 ± 5.25* 64.29 ± 0.95*
BrainGNN 64.07 ± 2.81* 67.16 ± 1.50* 61.93 ± 0.49* 60.73 ± 1.51* 62.34 ± 2.15*
AL-NEGAT 69.33 ± 3.56 68.60 ± 2.91* 65.27 ± 1.97* 63.14 ± 2.52* 66.88 ± 3.52
MVS-GCN 67.49 ± 2.91* 66.72 ± 3.05* 68.80 ± 1.60* 65.11 ± 4.13* 67.53 ± 1.11*
Hi-GCN 68.84 ± 2.15* 67.39 ± 2.76* 69.42 ± 2.74* 65.39 ± 3.33* 67.15 ± 3.90*
BrainNetCNN 68.33 ± 2.32* 66.71 ± 1.30* 64.81 ± 1.48* 64.01 ± 1.41 66.38 ± 2.90*
BrainNetTF 69.26 ± 3.09* 67.64 ± 2.39* 69.53 ± 0.83* 67.70 ± 2.92* 68.06 ± 1.53*
MAHGCN 68.34 ± 3.56 68.13 ± 1.88 68.22 ± 2.67* 67.91 ± 1.06* 68.67 ± 3.34*
Com-BrainTF 70.10 ± 2.65* 67.73 ± 2.45* 68.22 ± 0.96* 69.59 ± 0.73 69.23 ± 3.75*
@@
Table 3: Classification performance for ASD (ASD vs. NC) on ABIDE I and ABIDE II, and for ADHD (ADHD vs. NC) on ADHD-200, using the Brainnetome atlas. Results (shown as percentages) were derived from 10-fold cross-validation. Statistically significant differences from MHNet marked *. The best result in each category is highlighted in bold.
@@
ABIDE-I SVM 63.08 ± 3.19* 62.37 ± 3.82* 62.30 ± 4.83* 61.57 ± 2.84* 61.57 ± 3.35*
KNN 61.16 ± 3.39* 60.14 ± 2.65* 60.51 ± 2.56* 62.71 ± 1.35* 63.16 ± 1.17*
MLP 64.03 ± 3.89* 63.20 ± 4.62* 64.58 ± 3.25* 61.73 ± 2.80* 64.37 ± 2.46*
GCN 65.10 ± 6.59* 64.35 ± 1.62* 66.13 ± 2.34* 59.68 ± 6.52* 63.82 ± 1.27*
BrainGNN 69.94 ± 2.62* 68.68 ± 1.62* 66.28 ± 0.62* 65.68 ± 0.62* 66.23 ± 2.36*
AL-NEGAT 71.29 ± 3.19* 70.37 ± 2.54* 71.13 ± 2.17* 68.21 ± 2.39* 70.29 ± 3.10*
MVS-GCN 72.51 ± 2.12* 71.33 ± 2.92* 72.13 ± 2.42* 69.24 ± 3.85* 71.30 ± 1.39*
Hi-GCN 74.55 ± 3.23* 72.31 ± 2.64* 70.34 ± 2.53* 68.03 ± 2.45* 71.81 ± 3.59*
BrainNetCNN 72.59 ± 2.91* 71.56 ± 0.93* 72.21 ± 1.35* 67.59 ± 1.31* 72.24 ± 2.85
BrainNetTF 74.93 ± 2.37* 73.28 ± 1.97* 72.11 ± 1.27* 68.53 ± 2.71* 70.89 ± 1.73*
MAHGCN 73.21 ± 3.02* 73.52 ± 1.63* 73.02 ± 2.14* 71.55 ± 1.38 72.57 ± 3.03
Com-BrainTF 74.12 ± 1.72* 73.58 ± 1.81* 74.31 ± 0.51* 70.64 ± 0.38* 74.91 ± 3.41
@@
ABIDE-II SVM 61.43 ± 2.23* 60.15 ± 2.37* 61.25 ± 3.74* 59.44 ± 1.06* 60.57 ± 4.25*
KNN 62.37 ± 1.09* 61.24 ± 1.76* 63.83 ± 1.67* 60.82 ± 2.15* 62.07 ± 2.37*
MLP 62.83 ± 2.42* 62.02 ± 3.00* 63.15 ± 2.61* 60.54 ± 2.41* 63.15 ± 1.98*
GCN 63.80 ± 2.76* 62.85 ± 2.03* 64.96 ± 3.62* 58.17 ± 3.87* 62.65 ± 1.84*
BrainGNN 69.37 ± 3.27* 66.34 ± 2.35* 67.71 ± 2.49* 65.28 ± 3.74* 67.77 ± 2.75*
AL-NEGAT 70.14 ± 0.80* 68.11 ± 0.57* 70.95 ± 3.41* 68.95 ± 3.22* 69.11 ± 3.55*
MVS-GCN 69.24 ± 3.93* 67.01 ± 2.30* 69.97 ± 2.12* 67.80 ± 3.23* 68.92 ± 3.91*
Hi-GCN 70.30 ± 2.74* 69.94 ± 1.00* 70.17 ± 3.81* 68.64 ± 2.33* 69.86 ± 1.95*
BrainNetCNN 69.32 ± 1.43* 67.13 ± 3.21* 68.02 ± 2.10* 65.14 ± 2.49* 66.98 ± 0.57
BrainNetTF 71.42 ± 2.66* 70.89 ± 2.64* 71.97 ± 1.66* 69.03 ± 2.80* 71.76 ± 2.89*
MAHGCN 72.17 ± 1.76* 71.27 ± 2.03* 72.56 ± 2.94* 69.89 ± 0.71 71.33 ± 2.83
Com-BrainTF 72.73 ± 2.85* 71.18 ± 1.24* 72.92 ± 0.95* 69.41 ± 1.60* 71.35 ± 1.77
@@
ADHD-200 SVM 61.15 ± 2.21* 60.20 ± 3.00* 60.30 ± 4.31* 59.65 ± 2.04* 59.83 ± 3.14*
KNN 59.09 ± 2.71* 58.38 ± 2.42* 58.09 ± 1.61* 60.18 ± 0.78* 61.47 ± 0.70*
MLP 61.75 ± 3.55* 61.30 ± 3.82* 62.63 ± 2.74* 59.26 ± 2.19* 63.14 ± 2.34*
GCN 60.84 ± 5.93* 63.19 ± 0.97* 63.79 ± 1.68* 57.42 ± 5.57* 61.43 ± 3.56*
BrainGNN 64.66 ± 2.62* 62.78 ± 1.62* 62.33 ± 0.62* 61.21 ± 0.62* 63.00 ± 2.36*
AL-NEGAT 64.03 ± 3.19* 63.21 ± 2.54* 62.79 ± 2.17* 63.95 ± 2.39* 63.90 ± 3.10*
MVS-GCN 66.96 ± 2.12* 64.04 ± 2.92* 66.33 ± 2.42* 63.98 ± 3.85* 65.05 ± 1.39*
Hi-GCN 67.49 ± 3.23* 66.97 ± 2.64* 65.84 ± 2.53* 64.14 ± 2.45* 66.08 ± 3.59*
BrainNetCNN 69.08 ± 2.91* 68.31 ± 0.93* 69.42 ± 1.35* 65.64 ± 1.31* 68.08 ± 2.85
BrainNetTF 68.90 ± 2.37* 68.26 ± 1.97* 68.29 ± 1.27* 66.03 ± 2.71* 67.37 ± 1.73*
MAHGCN 69.56 ± 3.02* 67.76 ± 1.63* 68.88 ± 2.14* 68.33 ± 1.38 68.47 ± 3.03
Com-BrainTF 70.12 ± 1.72* 68.14 ± 1.81* 70.99 ± 0.51* 67.05 ± 0.38* 69.25 ± 3.41
@@ 5.2.1 Influence of Cutoff Threshold
For HGNN, we use the obtained RV coefficient matrix to construct the adjacency matrix of the brain hierarchical network of each data set, but the RV coefficient matrix is not a sparse binary matrix. In order to achieve the best classification effect, threshold processing is required for the matrix. Specifically, if the RV coefficient between brain area $i$ and brain area $j$ is less than the cutoff threshold $\gamma$, the value at $(i,j)$ of the RV coefficient matrix is set to 0. As shown in Figure 3, in order to select a suitable cutoff threshold, we plotted the relationship curve between the percentage of retained edges and the cutoff threshold, and selected the inflection point of the curve as the cutoff point. The cutoff threshold at this point is used to process the RV coefficient matrix into a sparse binary adjacency matrix.